OpenStudy (anonymous):

how to solve this integral ? i tried by taking x^3=z

OpenStudy (anonymous):

\[\int\limits \frac{ dx }{ x(1+4x^3+3x^6 )}\]

OpenStudy (anonymous):

Notice that \(1+4x^3+3x^6=0\) when \(x=-1\). This means you can factor the 6th degree polynomial. Using long division, you get